Can anyone answer this question in calculus?
for Find the particular soln. the folwing D.E. 1) 9-44-129=2x-x+3 ans: gp =- 1 x + 2 x² - 1/3 X - 5 / 44 ans - x 6
Find the the particular soln. for folwing D.E. 2. g - g -2y = sin 2x + x ansi. Yp- of the cos 2x - 3 2 sin 2x + 1 *
Find the the particular soln. for folwing D.E. 3. 9-9 = x 3 x aus: yerli - Base
Find the particular soln. for the folwing D.E. 43 g -49=25in 3x-3 6053x aus: Yp=-2 sin 3x + 3 Cos 3
Find the particular soln. for the folwing D.E. g +4g=sin 3x-5 cosx 5
Find for the the particular soln. folwing D.E. 6 g -99= 6*²-7 sin3x
Find the particular soln. for the folwing D.E. 3 g -49-124 = 3 et + siu zx 4X
Answer #1

03-19 -12-27 “ = 1 9-12 l 2^2-x 3 + 2xx-x+3) <p 9-02 J+ = D -11 z , 2013-90 (39) 1999) [* - J+ ]- + f/ [*(= ) - -)(--)]0 g-yl-2y = sinza test sinaat _ 394- 3- 2 e 5 --0-2) Sinax + 3 os sinax & I - a - 2-I (D+o) Sin2x + I ( +6) (-6) -_CD-6) Si② y-y = a ex 323 x = [10+3)²-1 3 *20*4872X30- 2) e 3x 02 +60 + =4y = 2 sin 32 - 368 3x 3p = Oku 2 sinux - L_300834 D -4) (06 E y + 4y = sinan -50.8% Yp= t sinx-s t cosa (02-4) (02-4) E ou sin 32 - _5. Cosa yp to sinox + scosse y = - sin3x + comm] ©25-4X5-12 -4-40-12 10- _ 34 +-40-16 3 y - yy-12y= 3 nt sin axta eyn yp intuneset-L 02-40-12) sinax tine! (03-40-12) (02-40-